OpenStudy (anonymous):

Jonah deposited $500 in a savings account, which pays 2.5 percent interest. If he keeps it in the account for 3 and a half years, how much interest will he earn?

OpenStudy (anonymous):

Is it compounding interest and annual interest?

OpenStudy (anonymous):

Year 0 to 1: 1.025 * 500 = 512.5 Year 1 to 2: 1.025 * 512.5 = 525.3125 Year 2 to 3: 1.025 * 525.31 = 538.4453125 Year 3 to 3.5: 1.0125 * 538.44 = 545.17587890625 Then I rounded it to $545.18